(5√x-2) - 3 = (√19x-29)
I've been able to solve problems with two rationals, but the 5 added before the first one is causing me problems. step-by-step solutions please!

Answer by nyc_function(2741)

5 √[x - 2]) - 3 = √[19x - 29] square both side

25(x-2) – 30√(x-2) + 9 = 19 x – 29 add by -19x +30√(x-2) +29

6x-12 =30√(x-2) divide by 6

X -2 = 5√ (x-2) square both side

X^2 -4x + 4 =25x – 50

X^2 -29x +54=0

(x-27)(x-2) =0 → x= 27

We reject x = 2 because it does not satisfy the original radical equation.

The only answer is x = 27.